Solve the diffrential equation sin inverse dy/dx=x+y

Grade:12

1 Answers

Aditya Gupta
5 years ago
put z=x+y 
so that dz/dx= 1+dy/dx=1+z
or dz/1+z = dx
or ln|1+z|= x+c
or 1+z= ke^x
or 1+x+y= ke^x, k being an arbitrary constant